OpenStudy (anonymous):

I need a formula for solving this. Not the answer a formula. A basketball team has ten players. 5 are gaurds, 3 are centers, 2 are forwards. If there has to be two gaurds two centers and one forward how many possible combinations are there?

OpenStudy (anonymous):

\[\sqrt{?} \] Also what does this mean?

OpenStudy (anonymous):

second question first. it means the positive square root

OpenStudy (anonymous):

so for example \[\sqrt{25}=5\]

OpenStudy (anonymous):

now for first question. you need 2 guards. the number of way to pick two people from a set of 5 is \[\frac{5\times 4}{2}=10\] this is sometimes written as 5C2 or \[\dbinom{5}{2}\]

OpenStudy (anonymous):

there are 3 ways to pick 2 centers from a set of 3. just like saying "how many ways can i exclude 1 of them" and there are obviously 3 ways. if you want to write 3C2=3 that is fine too

OpenStudy (anonymous):

obviously 2 different ways to choose one forward from the 2 you have. so your "final answer" by the counting principle is \[10\times 3\times 2=60\]

OpenStudy (anonymous):

Thanks :)